Grade Upto college level Engineering Entrance Exams

The tension in the sonometer wire is increased by 21%. It fundamental frequency (a) Increases by 21% (b) Increases by 10% (c) Increases by 5% (d) Increases by 7%

sumit kumar
12 Years agoGrade Upto college level

1 Answer

Vikas TU
6 Years ago
fundamental frequency
________
f=1/2l√i/m
length of the wire
f¹/f. =(t¹/t)½
hence
f¢= f+6
t¢=t+0.44t
so ,
______________
(f+6)/f = √ 1.44t/t
f=30 Hertz

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago